# and quantum information science

### v1

@@ 8.2 Holographic path-integral optimization
The optimization process described above starts with a CFT path integral in a flat geometry, and effectively places it on a particular hyperbolic metric. After an appropriate identification of the Euclidean time with a radial coordinate, this metric can be interpreted as a slice of AdS spacetime. It is then natural to ask whether this procedure, when applied to holographic CFTs, could have a dual geometric interpretation within the AdS/CFT correspondence. If successful, this embedding in holography could shed new light on the interpretation of the gravity dual of the Liouville action in terms of path-integral complexity of a CFT state.
With this motivation, references [368, 369] first proposed a holographic dual of the path-integral optimization procedure. Their construction was very closely related to the framework of Anti-de Sitter/Boundary Conformal Field Theory (AdS/BCFT) [336]. The main idea was to propose as the geomtrical dual the maximization of Hartle-Hawking (HH) wave functions in gravity. We consider a boundary cut-off surface $\Sigma$ located at $z=\epsilon$ (where $z$ is the radial coordinate of the AdS spacetime), and a generic spacelike surface $Q$ parametrized by $z=f(\tau)$. Referring to Fig. 14, the HH wave functions were computed using Eq. (130) inside the shaded region $M$ bounded by $\Sigma$ and $Q$. The wave functions need to be maximized with respect to the induced metric $h_{ab}$ on the time-slice $Q$, described by the metric (with an appropriate choice of gauge)
@@
Then, to the leading order in $1/G_{N}$, the HH wave function can be semiclassically computed as the exponent of the gravitational action. The latter is composed by the Einstein-Hilbert bulk term, plus appropriate boundary contributions: the Gibbons-Hawking term, and a codimension-one brane term with tension $T$. As a result, the maximization in gravity is equivalent to imposing the Neumann boundary conditions on $Q$ (indeed in complete analogy with AdS/BCFT)
@@
where $K$ is the trace of extrinsic curvature $K_{ab}$ on $Q$. Interestingly, performing the optimization of HH wave functions in AdS3 spacetimes as well in the BTZ black hole background, reproduces metrics identical to those in the CFT optimization (131), after the following identifications:
@@
One key insight from this construction was the UV completion of the Liouville action, which could be interpreted as the leading $1/\epsilon$ contribution to path-integral complexity (see also the discussion in [367] and its summary in the preceding section). Specifically, [368, 369] argued that the gravity action in region $M$ (see Fig. 14), supplemented with appropriate boundary and Hayward corner terms, provided a finite-cutoff completion of the Liouville action. As a consistency check, taking the "boundary" limit of the gravity action reproduced both the Liouville action (133), and its higher-dimensional generalization [29].
A closely related setup was later explored in [370] as an approach to holographic tensor networks in holographic CFTs. It was demonstrated that the CFT on the slices $Q$ could be explicitly constructed using $T\bar{T}$ deformation with a coupling that depended on Euclidean time [^31] This deformation later became the starting point of the “Cauchy-slice holography” [371]., and could be described within finite cut-off holography [372]. Parallel works [31, 33] also argued for $T\bar{T}$-like interpretation of path integral complexity in holographic CFTs.
The key distinction between references [368, 369] and the work [370] was the way in which partition functions of holographic CFTs deformed by $T\bar{T}$ were computed. Namely, in the $T\bar{T}$ context, the on-shell action was evaluated in the bulk region from the interior up to $Q$; whereas in the holographic path-integral optimization context, the action was evaluated in the region $M$ attached to the boundary. In both cases, HH wave functions give rise to identical, constant mean curvature slices of AdS3 spacetime. Interestingly, slices of AdS spacetime with constant extrinsic curvature also appear naturally in the context of tensor networks. Reference [338] used the same kind of slices to match the volume of the ERB in three-dimensional black hole geometries with a boundary notion of Nielsen complexity.
Concluding, path integral optimization has a great advantage: it defines complexity as a partition function, which is central for a precise holographic map between quantum and gravitational observables. It would be interesting to further push this notion of complexity to find a precise holographic map to geometric observables in AdS spacetime.
@@ 9 Quantum complexity in quantum information theory and many-body physics
Quantum information (QI) theory and complexity inform each other in many ways. We focus on three. First, QI theory shows that complexity governs a resource consumed in operational tasks (Sec. 9.1). Second, complexity distinguishes topological phases (Sec. 9.2). Third, a QI-theoretic quantity—entanglement entropy—bounds complexity under certain conditions (Sec. 9.3). [^32] We already discussed another connection between complexity and entanglement in the context of Nielsen’s complexity in section 5.4. There, we saw that very special definitions of complexity are completely fixed by the entanglement and Rényi entropies. Finally, the output of probability distributions also serves as a tool to bound complexity (Sec. 9.4). We chose these topics because holographers called for the resource theory, entropy is the workhorse of QI theory, and the topological application predated holographic applications of complexity. Other QI-complexity intersections lie outside this review’s scope. Examples include the bounding of complexity with the quantum Wasserstein distance [373], the relationship between complexity and quantum circuits’ expressivity [374], and applications of complexity to quantum machine learning [375].
@@ 9.1 Operationalism and resource theory
Brown and Susskind suggested that complexity controls a resource useful in operational tasks [44]. Let $\rho$ denote an arbitrary quantum state of a system $S$; $\mathcal{C}(\rho)$, the complexity of $\rho$; and $\mathcal{C}_{\rm max}$, the greatest possible complexity of $S$. Define the uncomplexity of $\rho$ as the gap $\mathcal{C}_{\rm max}-\mathcal{C}(\rho)$. Uncomplexity serves as a resource in quantum computation: consider running a quantum computation on an $n$-qubit register. The optimal input state is the zero-complexity state $|0\rangle^{\otimes n}$. It serves analogously to clean scrap paper on which we write scratchwork when performing classical computations.
Brown and Susskind conjectured that one can define a resource theory for quantum uncomplexity [44]. A resource theory is a simple, information-theoretic model for any context in which agents are restricted to perform only certain operations. For example, in the resource theory of entanglement, agents can perform only local operations and classical communications (LOCC). The allowed operations are called free, being modeled as costing nothing. Anything not free is a resource; it may be consumed to enable an operational task. In the entanglement theory, for instance, entangled states are resources. Agents may consume a maximally entangled pair of qubits to teleport quantum information, using LOCC.
Yunger Halpern et al. defined a resource theory of quantum uncomplexity in [376]. For simplicity, suppose that an agent in the resource theory has $n$ qubits (the setting can be generalized). The free operations are called fuzzy operations: the agent can attempt to perform any two-qubit gate $U$ (the setting can be generalized to $k$-qubit gates, for any $k\leq n$, and to gates in a particular set). However, noise corrupts every gate implementation. Hence the gate $\tilde{U}$ realized is chosen randomly from an $\epsilon$-ball (a small open set) about $U$. $\tilde{U}$ is called a fuzzy gate. The compositions of all fuzzy gates are the fuzzy operations. Every agent will have a personal tolerance for how fuzzy, or unknown, the state can become. This tolerance limits how many fuzzy gates the agent is willing to perform—and so curtails the complexities of the operations performed on the system.
Resource theories can be used to formulate operational tasks and to quantify the tasks’ optimal efficiencies. An example is uncomplexity extraction [376]: let $\delta\geq 0$ denote an error tolerance; and $\rho$, an arbitrary $n$-qubit state. One extracts uncomplexity from $\rho$ by distilling a state $\delta$-close to $|0\rangle^{\otimes k}$ in trace distance. How large a $k$ is achievable? The answer depends on the complexity entropy.
@@
These results offer hope that complexity entropies might quantify holographic tasks’ operational efficiencies. For example, consider an agent Alice falling into a black hole [44]. A firewall is expected to consume her if the horizon is dual to a CFT whose complexity is decreasing. Tossing a qubit into the black hole doubles the CFT state’s uncomplexity, enabling the complexity to return to decreasing—and saving Alice. This story might be generalized beyond qubits with help from $H_{\rm c}^{r,\eta}$.
@@ 9.2 Circuit complexity and topological order
Up until the early 80s it was thought that all phases of matter fit into Landau’s framework of symmetry breaking [381, 382]. Roughly, Landau’s theory asserts that a phase transition can only occur when a symmetry is broken, like the spontaneous change from a continuous symmetry in a chaotic/liquid phase to a discrete symmetry in a solid phase. Here, every phase is described by a pair of symmetry groups $G_{H}$ and $G_{\psi}$, where $G_{H}$ are all symmetry operations of the Hamiltonian and $G_{\psi}\subseteq G_{H}$ is the symmetry group of the ground state $\psi$. A phase transition then occurs if the symmetry group $G_{\psi}$ changes under an order parameter (such as the strength of a magnetic field).
@@
Starting with a plethora of inequivalent chiral spin states all satisfying the same symmetries, it was discovered that quantum Hamiltonians exhibit an even richer theory of phase transitions. For an extensive review about topological phases, we refer to Wen [383]. Since then an entire “zoo of topological phases” was discovered [384]. Topological phases refer to gapped ground states that show vast differences in their entanglement structure but are otherwise seemingly featureless (and therefore indistinguishable by symmetry groups).
@@
In the following we will show how quantum circuit complexity is related to the concept of topological order. States exhibiting topological order are gapped ground states of local Hamiltonians that cannot be reduced to a product state continuously without the gap closing. In particular, they are in a topological phase distinct from all product state, also called the trivial phase. More precisely, consider a gapped local Hamiltonian $H_{0}$ with a unique ground state $|\psi\rangle$. $|\psi\rangle$ can always be transformed into a product state via a continuous path consisting of local Hamiltonians $H(s)$ for $s\in[0,1]$ with $H(0)=H_{0}$: For $H_{0}$ acting on $n$ qubits we simply define $H(s)=(1-s)H_{0}+s\sum_{i=1}^{n}Z_{i}$, where $Z_{i}$ denotes the Pauli $Z$ matrix acting on the $i$th qubit/spin. Clearly, the unique ground state of $H(1)$ is $|0^{n}\rangle$ and $H(1)$ is gapped. However, it is very likely that at least one of the $H(s)$ has a degenerate ground state, and thus closes the gap. A state is in the trivial phase if such a continuous path exists such that the gap above the unique ground state $|\psi(s)\rangle$ is uniformly lower bounded by a positive constant along the path. A state is topologically ordered if no such path exists. A continuous path of Hamiltonians with a uniformly lower bounded spectral gap is called adiabatic.
@@
How does this relate to quantum circuit complexity? It was observed in Ref. [385] that the above definition of topological order in terms of the absence of adiabatic paths is equivalent to superconstant lower bounds on the quantum circuit complexity. It turns out that that the existence of an adiabatic path is equivalent to the existence of a local unitary transformation [385]:
@@ 9.3 Bounding complexity growth with entanglement
Quantum circuit complexity of a state quantifies the elementary resources necessary to prepare said state from a product state. Therefore, complexity also quantifies the resources necessary to transform the state back into a product state. Clearly, complexity can be viewed as a measure of entanglement, which is ultimately the failure of a state to be product. As such, we can compare it to the entanglement entropy, arguably the most prominent measure of entanglement. A high entanglement entropy along a bipartite cut does not imply a high circuit complexity: the maximally entangled state between two systems each consisting of $n$ qubits can be prepared on a quantum computer with a constant depth quantum circuits. Even if we consider the maximum entanglement entropy over all bipartite cuts of the qubits, we find that it saturates trivially at a maximum value of $n$, whereas the circuit complexity grows for an exponentially long time eventually reaching values of $\Omega(2^{n}/n)$ (compare Section 4.3). On the other hand, we can easily show the existence of quantum circuits with superpolynomial circuit complexity but no entanglement between some parts of the system. E.g., simply apply a Haar random unitary to $\log^{1+\varepsilon}(n)$ many qubits and leave all other qubits untouched. So if these two notions are not equivalent, what can we say about their relationship?
Entanglement growth implies early-time bounds on complexity during chaotic evolutions. More precisely, it turns out that entanglement does imply non-trivial circuit lower bounds, but entanglement needs to be present along any bipartite cut. The reason behind this is that the Schmidt rank along any bipartite cut can only be increased by a constant factor. In particular, if a state $|\psi\rangle$ has a Schmidt rank of $\mathrm{SR}_{i}$ for a bipartite cut of all qubits into two sets $A=\{1,\ldots,i\}$ and $B=\{i+1,\ldots,n\}$, then we know that at least $c\log(S_{i})$ gates (each acting on a qubit in $A$ and a qubit in $B$) are required to prepare $|\psi\rangle$. Suppose that we only want to lower bound the size of any geometrically local circuit that approximates the state $|\psi\rangle$. However, a geometrically local circuit can only apply gates to neighbouring qubits $(i,i+1)$. Then, any gate that acts on $A$ and $B$ necessarily acts on $(i,i+1)$. We can thus apply the above argument for all $n$ Schmidt ranks $\mathrm{SR}_{i}$ without overcounting and find:
@@
A version of this argument was applied in Ref. [106] to lower bound the circuit cost by the entanglement cost, which is closer to the more natural entanglement entropy than the Schmidt rank we used for the argument above. This bound can never produce a superpolynomial lower bound on the circuit complexity $\log(\mathrm{SR})=O(n)$.
Finally, let us mention a relation between circuit complexity and the so-called embezzlement of entanglement. The latter terminology refers to a process in which entanglement is approximately extracted from a resource system by using local unitary operators, with no detectable change to the resource’s state  [386]. References [387, 388] argued that any state of a relativistic QFT could serve as a resource for the process, where the error of the approximation can be made arbitrarily small. Recently, the author of [389] showed that circuit complexity is lower-bounded by a quantity proportional to the inverse of the error. In other words, circuit complexity acts as an obstruction to perfect embezzlement.

### v2

@@ 8.2 Holographic path-integral optimization
The optimization process described above starts with a CFT path integral in a flat geometry, and effectively places it on a particular hyperbolic metric. After an appropriate identification of the Euclidean time with a radial coordinate, this metric can be interpreted as a slice of AdS spacetime. It is then natural to ask whether this procedure, when applied to holographic CFTs, could have a dual geometric interpretation within the AdS/CFT correspondence. If successful, the embedding in holography could shed new light on the interpretation of the gravity dual of the Liouville action in terms of path-integral complexity of a CFT state.
With this motivation, references [421, 422] proposed a holographic dual of the path-integral optimization procedure. Their construction was closely related to the framework of Anti-de Sitter/Boundary Conformal Field Theory (AdS/BCFT) [387] in which one considers only a portion of the holographic geometry bounded by an end-of-the-world surface.
More precisely, consider a computation of the Euclidean Hartle-Hawking wavefunction in a portion $M$ of AdS spacetime on Fig. 14. The bulk region is bounded by the boundary cut-off surface $\Sigma$ located at $z=\epsilon$ (where $z$ is the radial coordinate of the AdS), and a generic spacelike surface $Q$ parametrized by $z=f(\tau)$ (homogeneous in the $x$ direction). The induced metric on $Q$ can be put into a Weyl-flat form by introducing an appropriate coordinate $\omega$ (i.e. a choice of gauge)
@@
The HH wave function is then a functional of the induced metric defined by path integral over all bulk metrics with this boundary condition on $Q$ (as well as the implicit boundary condition on $\Sigma$)
$\Psi_{HH}[\phi]=\int[Dg_{\mu\nu}]e^{-I_{G}[g]}\delta(g_{ab}|_{Q}-e^{2\phi}\delta_{ab})\,,$ (150)
where $I_{G}[g]$ stands for the Einstein-Hilbert action with the boundary Gibbons-Hawking term. Computing this path integral is in general a formidable task but, to the leading order in $1/G_{N}$, the semiclassical HH wave function becomes the exponent of (minus) the on-shell gravitational action itself. Mimicking the AdS/BCFT construction, we can also supplement the Gibbons-Hawking term by a constant tension parameter $T$ on $Q$ modeling certain types of bulk matter (and boundary entropy in AdS/BCFT).
Now, the main observation of [421, 422] was that a maximization of the semiclassical HH wave function with respect to the choice of $\phi(\omega)$ is equivalent to imposing the Neumann boundary conditions (in analogy with AdS/BCFT) on $Q$
@@
where $K$ is the trace of extrinsic curvature $K_{ab}$ on $Q$. Moreover, performing these steps for HH wave functions in Poincare, global and (static) black hole spacetimes, yields “maximizing" metrics (149) that match those in the CFT optimization (144), after the following identifications:
@@
This agreement was conjectured (interpreted) as the correspondence between the maximization of HH wave functions in AdS and the path integral optimization in holographic CFTs. It is also interesting to comment that parameter $\mu$ in the Liouville action effectively quantifies the amount of optimization, ranging from the unoptimized original metric for $\mu=0$ and the optimal for $\mu=1$. This way, tension $T$ can be seen as its holographic counterpart as it parametrizes a family of surfaces $Q$ starting from the boundary surface $\Sigma$ for $T=-(d-1)$ up to surface $\tau=0$ for $T=0$ (see [421] for detailed discussion).
The second key insight from this construction was a UV completion of the Liouville action, which could be interpreted as specifiying the leading $1/\epsilon$ contribution to path-integral complexity (see the discussion in [420]). Specifically, [421, 422] argued that the gravity action in region $M$ (see Fig. 14), supplemented with appropriate boundary and Hayward corner terms, provides a finite-cutoff completion of the Liouville action. As a consistency check, taking the "boundary" limit of the gravity action reproduced both the Liouville action (146), and its higher-dimensional generalization [39].
A closely related setup was later explored in [423] as an approach to tensor networks in holographic CFTs. The authors demonstrated that the CFT on the slices $Q$ above could be explicitly constructed using $T\bar{T}$ deformation with a coupling that depends on Euclidean time, [^41] This deformation later became the starting point of the “Cauchy-slice holography” [424]. and could be described within finite cut-off holography [425]. Parallel works [41, 43] also argued for $T\bar{T}$-like interpretation of path integral complexity in holographic CFTs. The key distinction between [421, 422] and [423] was the way in which the partition functions of holographic CFTs deformed by $T\bar{T}$ were computed. Namely, in the $T\bar{T}$ context, the on-shell action was evaluated in the bulk region from the interior up to $Q$; whereas in the holographic path-integral optimization context, the action was evaluated in the region $M$ attached to the boundary. Nevertheless, as functionals of the induced metric on $Q$ (i.e. the Weyl factor $\phi(\omega)$), gravitational actions in both computations are the same; therefore, their maximization yields the same metrics $\phi(\omega)$. In other words, their maximization selects a particular slice $z=f(\tau)$ of asymptotically, locally $\mathrm{AdS}$ spacetimes that has constant extrinsic curvature $K$, the so-called constant-mean-curvature (CMC) slices.
Interestingly, slices of AdS spacetime with constant extrinsic curvature also appear naturally in the context of bulk and boundary symplectic forms [388]. Reference [48] proposed the same kind of slices to match the volume of the Einstein-Rosen Bridge in three-dimensional black hole geometries with a boundary notion of Nielsen complexity.
In summary, path integral optimization defines complexity via a partition function, a formulation which lends itself naturally to a description via the holographic map between CFT and gravitational observables. It would be interesting to make this map more precise.
@@ 9 Quantum complexity in quantum information theory and many-body physics
Quantum information (QI) theory and complexity inform each other in many ways. We focus on three. First, QI theory shows that complexity governs a resource consumed in operational tasks (Sec. 9.1). Second, complexity distinguishes topological phases (Sec. 9.2). Third, a QI-theoretic quantity—entanglement entropy—bounds complexity under certain conditions (Sec. 9.3). [^42] We already discussed another connection between complexity and entanglement in the context of Nielsen’s complexity in section 5.5. There, we saw that very special definitions of complexity are completely fixed by the entanglement and Rényi entropies. Finally, the probability distribution over bitstrings $|x\rangle,x\in\{0,1\}^{n}$ induced by measuring a state also serves as a tool for bounding complexity (Sec. 9.4). Other QI-complexity intersections that we do not discuss extensively below include the bounding of complexity in terms of the quantum Wasserstein distance [426], the relationship between complexity and the expressivity of quantum circuits [427], and applications of complexity to quantum machine learning [428].
@@ 9.1 Operationalism and resource theory
The authors of [55] suggested that complexity controls a resource useful in operational tasks. Let $\rho$ denote an arbitrary quantum state of a system $S$ and let $\mathcal{C}_{0}(\rho)$ denote the exact circuit complexity of $\rho$: the least number of basic operations required to prepare $\rho$. Let $\mathcal{C}_{\rm max}$ denote the greatest possible complexity of any state on $S$. For an $N$-qubit system, $\mathcal{C}_{\rm max}\sim 2^{N}\,$. Define the uncomplexity of $\rho$ as the gap $\mathcal{C}_{\rm max}-\mathcal{C}_{0}(\rho)$. Uncomplexity serves as a resource in quantum computation. For example, consider running a quantum computation on an $n$-qubit register. The natural starting point is the zero-complexity state $|0\rangle^{\otimes n}$. It serves analogously to clean scrap paper on which we write scratchwork when performing classical computations.
Brown and Susskind conjectured that one can define a resource theory for quantum uncomplexity [55]. A resource theory models scenarios in which agents are restricted to perform only certain operations. For example, in the resource theory of entanglement, agents can perform only local operations and classical communications (LOCC). The allowed operations are called free, and are modeled as costing nothing. Anything not free is a resource; it may be consumed to enable an operational task. In the entanglement theory, for instance, entangled states are resources. Agents may consume a maximally entangled pair of qubits to teleport quantum information, using LOCC.
A resource theory of quantum uncomplexity was defined in [429]. For simplicity, suppose that an agent in the resource theory has $n$ qubits (the setting can be generalized). The agent can attempt to perform any two-qubit gate $U$ (the setting can be generalized to $k$-qubit gates, for any $k\leq n$, and to gates in a particular set). However, noise corrupts every gate implementation. Hence the gate $\tilde{U}$ that is realized is chosen randomly from an $\epsilon$-ball (a small open set) about $U$. $\tilde{U}$ is called a fuzzy gate. The compositions of all fuzzy gates are the fuzzy operations. Every agent will have a tolerance for how fuzzy, or unknown, the state can become. This tolerance limits how many fuzzy gates the agent is willing to perform—and so curtails the complexities of the operations that will be performed on the system.
Resource theories can be used to formulate operational tasks and to quantify their optimal efficiencies. An example is uncomplexity extraction [429]: let $\delta\geq 0$ denote an error tolerance; and $\rho$, an arbitrary $n$-qubit state. One extracts uncomplexity from $\rho$ by distilling a state $\delta$-close to $|0\rangle^{\otimes k}$ in trace distance. How large a $k$ is achievable? The answer depends on the complexity entropy.
@@
These results offer hope that complexity entropies might quantify the operational efficiencies of holographic tasks. For example, consider an agent Alice falling into a black hole. A firewall is expected to consume her if the black hole state is dual to a CFT state whose complexity is not increasing [55]. [^43] A subtlety is that, when a black hole’s complexity stops increasing, classical general relativity is expected to stop describing the black hole accurately. Tossing a qubit into the black hole doubles the CFT state’s uncomplexity, enabling the complexity to return to increasing—and saving Alice. This story might be generalized beyond qubits with help from $H_{\rm c}^{r,\eta}$.
@@ 9.2 Circuit complexity and topological order
Until the early 80s it was thought that all phases of matter fit into Landau’s framework of symmetry breaking [434, 435]. Roughly, Landau’s theory asserts that a phase transition occurs when a symmetry is broken, like the spontaneous change from a continuous symmetry in a chaotic/liquid phase to a discrete symmetry in a solid phase. Here, every phase is described by a pair of symmetry groups $G_{H}$ and $G_{\psi}$, where $G_{H}$ are all symmetry operations of the Hamiltonian and $G_{\psi}\subseteq G_{H}$ is the symmetry group of the ground state $\psi$. A phase transition occurs if the symmetry group $G_{\psi}$ changes under an order parameter (such as the strength of a magnetic field).
@@
Starting with a plethora of inequivalent chiral spin states all satisfying the same symmetries, it was discovered that quantum Hamiltonians exhibit an even richer theory of phase transitions. For an extensive review about topological phases, see Wen [436]. Since then an entire “zoo of topological phases” has been discovered [437]. Topological phases refer to gapped ground states that show vast differences in their entanglement structure but are otherwise seemingly featureless (and therefore indistinguishable by symmetry groups).
@@
In the following we will show how quantum circuit complexity is related to the concept of topological order. States exhibiting topological order are gapped ground states of local Hamiltonians that cannot be reduced to a product state continuously without closing the gap. In particular, they are in a topological phase distinct from all product states, also called the trivial phase. More precisely, consider a gapped local Hamiltonian $H_{0}$ with a unique ground state $|\psi\rangle$. $|\psi\rangle$ can always be transformed into a product state via a continuous path consisting of local Hamiltonians $H(s)$ for $s\in[0,1]$ with $H(0)=H_{0}$. For $H_{0}$ acting on $n$ qubits, we simply define $H(s)=(1-s)H_{0}+s\sum_{i=1}^{n}Z_{i}$, where $Z_{i}$ denotes the Pauli $Z$ matrix acting on the $i$th qubit/spin. Clearly, the unique ground state of $H(1)$ is $|0^{n}\rangle$ and $H(1)$ is gapped. However, it is very likely that at least one of the $H(s)$ has a degenerate ground state, and thus closes the gap. A state is in the trivial phase if such a continuous path exists such that the gap above the unique ground state $|\psi(s)\rangle$ is uniformly lower bounded by a positive constant along the path. A state is topologically ordered if no such path exists. A continuous path of Hamiltonians with a uniformly lower bounded spectral gap is called adiabatic.
@@
How does this relate to quantum circuit complexity? It was observed in Ref. [438] that the above definition of topological order of a ground state $|\psi\rangle$ in terms of the absence of adiabatic paths is equivalent to a lower bound that grows in $n$ on the circuit depth for all circuits preparing $|\psi\rangle$ from a product state.
It turns out that that the existence of an adiabatic path is equivalent to the existence of a local unitary transformation [438]:
@@ 9.3 Bounding complexity growth with entanglement
Quantum circuit complexity of a state quantifies the elementary resources necessary to prepare it from a product state. Therefore, complexity also quantifies the resources necessary to transform the state back into a product state. Clearly, therefore, complexity can be viewed as a measure of entanglement, which is ultimately the failure of a state to be of product form. As such, we can compare it to the entanglement entropy, arguably the most prominent measure of entanglement. A high entanglement entropy along a bipartite cut does not imply a high circuit complexity: the maximally entangled state between two systems each consisting of $n$ qubits can be prepared on a quantum computer with a constant depth quantum circuit. Even if we consider the maximum entanglement entropy over all bipartite cuts of the qubits, we find that it saturates trivially at a maximum value of $n$, whereas the circuit complexity can be exponentially large, of order $\Omega(2^{n}/n)$ (compare Section 4.3). On the other hand, we can easily show the existence of quantum circuits with superpolynomial circuit complexity but no entanglement between some parts of the system. For example, simply apply a Haar random unitary to $\log^{1+\varepsilon}(n)$ many qubits and leave all the other qubits untouched. So the circuit complexity and the entanglement entropy are not equivalent; but what can we say about their relationship?
Entanglement growth implies early-time bounds on complexity during chaotic evolution. More precisely, it turns out that entanglement does imply non-trivial circuit lower bounds, but entanglement needs to be present along any bipartite cut [121]. The reason is that the Schmidt rank along any bipartite cut can only be increased by a constant factor. In particular, if a state $|\psi\rangle$ has a Schmidt rank of $\mathrm{SR}_{i}$ for a bipartite cut of all qubits into two sets $A=\{1,\ldots,i\}$ and $B=\{i+1,\ldots,n\}$, then we know that at least $c\log(\mathrm{SR}_{i})$ gates (each acting on a qubit in $A$ and a qubit in $B$) are required to prepare $|\psi\rangle$. Suppose that we only want to lower bound the size of any geometrically local circuit that approximates the state $|\psi\rangle$. A geometrically local circuit can only apply gates to neighbouring qubits $(i,i+1)$. Then, any gate that acts on $A$ and $B$ necessarily acts on $(i,i+1)$. We can thus apply the above argument for all $n$ Schmidt ranks $\mathrm{SR}_{i}$ without overcounting and find:
@@
At best, the left hand side is $\Omega(n^{2})$. This is because each Schmidt rank $\mathrm{SR}_{i}$ is trivially bounded by the state space dimension $2^{n}$. Therefore, it is impossible to prove exponential circuit complexity with this technique. Using similar tools, complexity was lower-bounded in one-dimensional models composed by noninteracting fermions [442]. Reference [443] argued that the area of Hartman-Maldacena surfaces (i.e., codimension-two minimal surfaces homologous to a boundary subregion [245]) lower-bounds the Schmidt rank of the unitary time evolution operator. In turn, the Schmidt rank lower-bounds the depth of any boundary circuit effecting such boundary unitary.
Next, let us mention a relation between circuit complexity and the so-called embezzlement of entanglement. The latter terminology refers to a process in which entanglement is extracted from a resource system by using local unitary operators, with approximately no detectable change to the resource state [444]. The embezzlement error estimates the amount by which this process changed the resource state. References [445, 446] argued that any state of a relativistic QFT could serve as a resource for the process, where the embezzlement error can be made arbitrarily small. Recently, the author of [447] showed that circuit complexity is lower-bounded by a quantity proportional to the inverse of the embezzlement error. In other words, circuit complexity acts as an obstruction to perfect embezzlement.
Entanglement, the subject of this subsection until now, does not suffice to enable quantum-computational speedups: classical computers can efficiently simulate even some highly entangled states’ evolutions. The power behind certain entangled states, which enables quantum-computational speedups, is called magic and nonstabilizerness. A possible relationship between magic and circuit complexity has been suggested recently.
To introduce the relationship, we first overview relevant background [448, 15]. Consider an $n$-qubit system. The Pauli group consists of $4^{n}$ Pauli strings, each a tensor product of $n$ Pauli operators and/or identity operators. The Clifford group is the Pauli group’s normalizer; each Clifford gate transforms every Pauli string into a Pauli string. The Hadamard, phase, and controlled-NOT gates span the Clifford group. The stabilizer states result from acting on $\lvert 0\rangle^{\otimes n}$ with Clifford gates. According to the Gottesman-Knill theorem, classical computers can efficiently simulate every quantum computation formed only from Clifford gates (such that the quantum computer always remains in the space of stabilizer states) [449]. Consequently, quantum-computational speedups stem from nonstabilizer states. The extra spice added by these states is called magic or nonstabilizerness [450]. This topic has recently advanced rapidly (e.g., [451, 452, 453, 454, 455, 456, 457, 458, 459, 460]). One can quantify magic in various ways, including with the stabilizer Rényi entropy, an entropic measure defined in terms of the Bell basis [453, 460]. The authors of [454] define a magic power that quantifies a unitary’s ability to behave unlike a Clifford gate—to evolve a Pauli string to a nontrivial linear combination of Pauli strings. The magic power lower-bounds the unitary’s circuit complexity [454]. The latter is quantified with Nielsen’s circuit complexity, particularly the cost function (25) evaluated at $p=q_{I}=1$.